OpenStudy (anonymous):

HELP! Which sum is equal to x^2+6x-5/x^2-25 ?

OpenStudy (anonymous):

do you have choices? this is not a sum, i think they want you to factor and cancel

OpenStudy (anonymous):

maybe it is \[\frac{x^2+6x+5}{x^2-25}\]?

OpenStudy (anonymous):

Or maybe using partial fraction composition? Sums of quotients?

OpenStudy (anonymous):

But yes, reads more like a problem where you factor and cancel.

OpenStudy (anonymous):

considering the other problems, i doubt that this is partial fractions

OpenStudy (anonymous):

but nothing cancels in any case, so i am not sure what to do with this one

OpenStudy (anonymous):

A. 1/x+5 + x+5/x-5 B. 1/x+5 + x/x-5 C. 1/x+5 + x^2/x+5 D. 1/x-5 + x/x+5

OpenStudy (anonymous):

that would be partial fractions

OpenStudy (anonymous):

Which one of those are equal to x^2+6x-5/x^2-25 ?

OpenStudy (anonymous):

i guess we have to add all of them and see what we get, because i am sure you do not know partial fraction decomposition unless i am wrong

OpenStudy (anonymous):

your right. ha

OpenStudy (anonymous):

oh good, I wont post that then (-:

OpenStudy (anonymous):

\[\frac{1}{x+5} + \frac{x+5}{x-5}\] \[=\frac{x-5+(x+5)(x+5)}{x^2-25}\] \[=\frac{x^2+10x+25+x-5}{x^2-25}\] \[=\frac{x^2+11x+20}{x^2-25}\] nope

OpenStudy (anonymous):

try the next one (it works)

OpenStudy (anonymous):

Okay!